# A Cascaded Random Access Quantum Memory

### v1

@@ I.3 Master equation simulation
The multimode cavity is fabricated from high-purity (99.9995%) Aluminum. Three rounds of chemical etching removed in total $\sim 200\mu$m surface using Transene Aluminum Etchant A.
For the transmon chip, we use a 430 $\mu$m thick C-plane HEMEX sapphire wafer annealed at $1200^{o}$ for 2 hours as the substrate. 150 nm thick Tantalum film is sputtered at $800^{o}$ as the ground plane. Large patterns, except Josephson junctions, are fabricated through photolithography with Heidelberg MLA 150 Direct Writer and fluorine-based dry-etching. The junction mask was fabricated with the Raith EBPG 5000+ E-Beam writer on a bi-layer resist (MMA EL11-950 PMMA A7). Transmon’s Josephson junctions are of the Manhattan type. The junction mask was metalized in a Plassys electron beam evaporator with double-angle evaporation. The wafer was then diced and lifted off and packaged inside the multimode flute cavity.
For the coupler chip, we use a 530 $\mu$m thick EFG C-plane sapphire wafer annealed at $1200^{o}$ for 2 hours as the substrate. 150 nm thick Niobium film is evaporated in Kurt J. Lesker E-beam Evaporator at room temperature as the ground plane. Large patterns, except Josephson junctions, are fabricated through photolithography with Heidelberg MLA 150 Direct Writer and Fluorine-based dry-etching. The junction mask was fabricated with Raith EBPG 5200+ E-Beam writer on a bi-layer resist (MMA EL13-950 PMMA A4). Coupler’s Josephson junctions are Dolan bridge type. The junction mask was metalized in a Plassys electron beam evaporator with double-angle evaporation. The wafer was then diced and lifted off. Each shorting line structure mentioned in Supplementary Section  J is scratched six times with an IPA-cleaned tweezer during packaging.
@@
Supplementary Fig. 16 shows the room and cryogenic temperature measurement chain. The device is mounted on the mixing chamber plate of a dilution refrigerator with a base temperature of 10 mK. All RF input signals are generated via the ZCU216 RFSoC board. Two DC sources (Yokogawa GS200) bias the DC flux of the coupler and the Josephson Parametric Amplifiers (JPA). One Signal Core SC5511A serves as the local oscillator (LO), split into two tones: one tone, after frequency up-conversion, functions as the readout pulse; the other is used for frequency down-conversion and is collected by the RFSoC ADC. A second Signal Core SC5511A provides the pump tone for the JPA, amplifying the qubit readout signal (labeled as Qout). The pulse sent to the qubit channel includes two components: the single-qubit drive and the qubit-buffer sideband drive, both synthesized directly through different RFSoC DAC channels. The sideband drive is further amplified with a ZHL-5W-422+ high-power amplifier at room temperature. Buffer and Storage modes are probed separately through a weakly coupled port machined at the multimode cavity; the signals are then combined and collected by the RFSoC ADC. Due to the RFSoC Nyquist frequency range, RF flux modulation is synthesized through two channels: low modulation frequency (0.1–1 GHz) and high modulation frequency (1–3 GHz). These two signals are combined at room temperature and then sent through the RF flux line, where they are combined with the DC flux at base temperature. The signal reaches the SMA connector, which is wire-bonded directly to the coupler chip. All RF lines have a DC block at the top of the fridge to avoid ground-loop, and all lines coming out of the device have Quantum Microwave eccosorbs placed as close as possible to the sample.

### v2

@@ Appendix I Random Read Fidelity
[…]
@@
Supplementary Figure 17 shows the room and cryogenic temperature measurement chain. The device is mounted on the mixing chamber plate of a dilution refrigerator with a base temperature of 10 mK. All RF input signals are generated via the ZCU216 RFSoC board. Two DC sources (Yokogawa GS200) bias the DC flux of the coupler and the Josephson Parametric Amplifiers (JPA). One Signal Core SC5511A serves as the local oscillator (LO), split into two tones: one tone, after frequency up-conversion, functions as the readout pulse; the other is used for frequency down-conversion and is collected by the RFSoC ADC. A second Signal Core SC5511A provides the pump tone for the JPA, amplifying the qubit readout signal (labeled as Qout). The pulse sent to the qubit channel includes two components: the single-qubit drive and the qubit-buffer sideband drive, both synthesized directly through different RFSoC DAC channels. The sideband drive is further amplified with a ZHL-5W-422+ high-power amplifier at room temperature. Buffer and Storage modes are probed separately through a weakly coupled port machined at the multimode cavity; the signals are then combined and collected by the RFSoC ADC. Due to the RFSoC Nyquist frequency range, RF flux modulation is synthesized through two channels: low modulation frequency (0.1–1 GHz) and high modulation frequency (1–3 GHz). These two signals are combined at room temperature and then sent through the RF flux line, where they are combined with the DC flux at base temperature. The signal reaches the SMA connector, which is wire-bonded directly to the coupler chip. All RF lines have a DC block at the top of the fridge to avoid ground-loop, and all lines coming out of the device have Quantum Microwave eccosorbs placed as close as possible to the sample.